OpenStudy (anonymous):

Y=ax^2+bx+c and vertex of the parabola is given (5,8) ; a=4 How many x intercepts does the parabola have?

OpenStudy (anonymous):

vertex is given by : \[(\frac{-b}{2a},\frac{-D}{4a})\],where D is discriminant of quadratic expression.

OpenStudy (anonymous):

\[(5,8)=(\frac{-b}{8},\frac{-D}{32})\] comparing x and y coordinate, \[\frac{-b}{8}=5\]\[b=-40\] and \[\frac{-D}{32}=8\] \[D=-256\] Now we know that if : D>0 , 2 distinct x-intercepts of parabola D=0 , 2 equal x-intercepts of parabola D<0 , no x-intercepts of parabola so here , parabola will have no x-intercepts